OpenStudy (anonymous):

Find f. f ''(t) = 5et + 3 sin t, f(0) = 0, f(π) = 0

OpenStudy (anonymous):

\[f''(t)=5e^t+3sin(t)\]?

OpenStudy (anonymous):

if so integrate these two seperately and it'll be easier \[f'(t)=\int{}{}{5e^t}dt+\int{}{}{3sin(t)dt}\]

OpenStudy (anonymous):

yepp. my bad ! it's ′′(t)=5e^t+3sin(t) but could youplz use Antiderivatives to solve it....?? cuz I havent studied about integrals.... ==''

OpenStudy (anonymous):

integrals = anti derivatives

OpenStudy (anonymous):

integrals are basically finding the area of something or finding volume and such but it uses the anti derivative to do this... basically mean't take the anti derivative of each

OpenStudy (anonymous):

do you understand how to use the anti derivative?

OpenStudy (anonymous):

i mean , for example, f′′(t)=5e^t+3sin(t) then F'(t)=5e^t-3cos(t)+C then F(x)=5e^t-3sin(t)+Cx+D

OpenStudy (anonymous):

that's correct

OpenStudy (anonymous):

so I just put 0 into F(x) then solve the C and D , right?

OpenStudy (anonymous):

now you just use your initial values to figure out what the constants are

OpenStudy (anonymous):

yep =]

OpenStudy (anonymous):

thanksssss :DDDDD

OpenStudy (anonymous):

0 will get you 0=5+0+0+d d=-5 and pi gets you 0=5e^{pi}+0+Cpi-5

OpenStudy (anonymous):

\[5e^{\pi}+C\pi\]=0 \[-5e^{\pi}=C\pi\] \[\frac{-5e^{\pi}}{\pi}=c\]

OpenStudy (anonymous):

you can try cleaning it up more if you'd like

OpenStudy (anonymous):

i'd just leave it

OpenStudy (anonymous):

hold on . I think my anti derivative of F(x) is not correct, dont I cuz if I put 0 and pi into the F(x) what I get are F(pi)=5e^pi-3sin(pi)+Cpi+D=0 F(0)=5e^0-3sin(0)+C*0+D=0------ 5-0+0+D=0---D=-5